I would appreciate any help on this problem, I've worked the word problem out from what examples were given in class but the answer I come out with doesn't seem correct when I check it. Please help with this problem:
One pipe can fill a pool in 8 HOURS. The drain can empty it in 14 HOURS. With the pipe on and the drain open, how long will it take the pool to be filled?

Answer by lheadle(1)

PIPE (fill)= 8 hours
PIPE (drain)= 14 hours
FILLEd = (-)
PIPE (fill)= RATE= 1/8 TIME= t WORK= t/8
PIPE (drain)=RATE= 1/14 TIME= t WORK= t/14
EQ: 1/8 - 1/14 = 1
: 112(1/8 - 1/14 = 1 )112
: 14t - 8t = 112
: 6t = 112
: 6t/6 = 112/6
: t = 18 4/6 or 18 2/3 HOURS